# A Goemans-Williamson type algorithm for identifying subcohorts in clinical trials

### v1

@@ 2.1 Underlying theoretical problem
Without loss of generality, we may assume that the input also specifies a vertex $u_{0}\in S_{U}$, since one can always iterate over all vertices in $U$. Note that, $S$ is large and $|S_{U}|\gg|S_{V}|$, together imply that the cut $(S_{U},V\setminus S)$ is large. Therefore, one may formulate the above as a MAX-CUT type integer program, where one associates a ±$\pm 1$ variable $z_{u}$ with each of the $n$ vertices in $G$, and wants to maximize the objective:
@@
    $\langle h,x_{u}\rangle\geq\theta(1-z_{u^{\prime}}z_{u_{0}})$ (3)
  • Sparsity: We want the designed test to use few genes, i.e., we want its coefficients to be sparse. Hence we bound the $\ell_{1}$ norm of its coefficients (see for example FR (13)) by an input parameter $s$, which is expressed as:
@@
It is worth noting that the formulated problem above is very similar to maximum margin classification using linear support vector machines. Thus, it is possible to generalize it to non-classifiers using the "kernel trick" (see for example BGV (92)). The only exception being the sparsity constraint above, which needs to be adapted for the chosen kernel. It may be possible to generalize the sparsity constraint for low degree polynomial kernels using standard compressed sensing methods (see for example FR (13)). However, such non-linear extensions are not the focus of this work.
@@
$\forall u^{\prime}\in V:$ $\ \langle h,x_{u^{\prime}}\rangle\geq\theta(1-\langle v_{u^{\prime}},v_{u_{0}}\rangle)$ (8)
@@
where $x_{u}$ are the embeddings in $\mathbb{R}^{d}$, $s$ is the sparsity parameter, and $\theta$ is the separation threshold parameter, all to be specified with the input. Algorithm 1 essentially maps the vector solution of the SDP to ±$\pm 1$ values, to obtain the subcohort (corresponding to $\{i\in U\cup V:\ z_{i}=1\}$).
@@
1:Input: Two sets of vertices $(U,V)$ labeled 00 and $1$ respectively, embedded in $\mathbb{R}^{d}$; vertex $u_{0}\in U$, sparsity parameter $s\in(0,\infty)$, and separation threshold $\theta\in\mathbb{R}$.
@@
4:Solve the SDP relaxation (Equation 5) to obtain the solution vectors $\{v_{i}\}_{n}$.
@@
6:$\forall i\in\{1,...,n\}:\ z_{i}\leftarrow\mathrm{sign}(\langle v_{i},\gamma_{i}\rangle)$, where $\gamma_{i}\sim\mathcal{N}(0,1)$.
@@
In order to show that the computed SDP solution above is not too far from the optimum, we prove the following approximation guarantee (see the supplement for details).
Theorem 2.2
Let $S^{*}=(S^{*}_{U},S^{*}_{V})$ denote the optimal subcohort, then Algorithm 1 computes a subcohort $S=(S_{U},S_{V})$, such that $|S_{U}|\geq\alpha|S^{*}_{U}|$, where the constant $\alpha\simeq 0.82$, when $|S_{U}^{*}|\ll|U|$.
Note that the optimum hyperplane separator $h$ remains unchanged after rounding though the input specified threshold $\theta$ is perturbed during rounding. However, we may bound the perturbation amount by using Grothendieck’s identity, i.e.,
@@
$\langle h,x_{u}\rangle$$\geq$ $\theta\left(1-\frac{2}{\pi}(z_{u^{\prime}}z_{u_{0}})\right)$ (11)
@@ 2.2 Application to genomics
  • In Subsection 2.2.1, we examine the quality of the identified subcohorts. A random test can be expected to generate subcohorts with less than 50% metastasis cases (as the number of non-metastasis patients is higher in the dataset). In Figure 1(a), we show that the linear tests generated by Algorithm 1 identify subcohorts where greater than 50% of the cases are metastatic (thus outperforming random). Moreover, if instead of using one hyperplane to identify a subcohort, we use two hyperplanes to identify a subcohort then the identified subcohorts have 60-70% of their cases as metastatic (see Figure 1(b)).
@@
  • In Subsection 2.2.2, we use the identified subcohorts to find statistically significant associations between epigenetic data and gene expressions. In SK (15), the authors identify about 40 genes to be of interest in breast cancers. Many of the genes are tumor suppressors, and methylation of the nucleotides in those genes leads to reduced expression and thus potentially cancer progression. Moreover, in BSW+ (13); OALH (06), the authors discuss the role of 50 nuclear receptors as drug targets. Therefore, it is natural to look for associations between significant changes in nuclear receptor expressions and significant changes in methylation levels of breast cancer related genes (over population baseline levels) – the latter can guide treatment drug targets from the former. We show that, even though at the entire dataset level such associations do not exist, but at the subcohort level they do exist (see Figure 3).
@@
Dataset: We use a subset of the METABRIC dataset CSC+ (12); PCR+ (16); RSS+ (19) [^7] The dataset is available at: https://www.cbioportal.org/study/summary?id=brca_metabric.that contains RNA microarray data from biopsy samples of approximately 1904 (911 with positive lymph nodes and 993 without positive lymph nodes) breast cancer patients, and methylation data for approximately 13K genes in about 1406 (669 with positive lymph nodes and 737 without positive lymph nodes). In this paper, we (crudely) denote patients with one or more positive lymph nodes upon biopsy as metastasis cases, while the rest as non-metastasis cases.
2.2.1 Predicting subcohorts in breast cancer metastasis
In order to generate Figure 1 (the main result of this subsection), Algorithm 1 takes as input the nuclear receptor expressions of a subset of 50 metastatic and 50 non-metastatic cases from the full dataset, the sparsity parameter $s$, a threshold $\theta$, and an arbitrary metastasis patient that is assumed to belong to the subcohort to be identified. The algorithm outputs a test $h$, where $h$ is just a tuple of weights for the nuclear receptors. On a separate hold-out set of a 1000 patients (500 with metastasis and 500 without), the test simply calculates the weighted average value of the given nuclear receptor expression levels for each patient in the hold-out, and if this value is below the threshold then the patient is added to the subcohort, otherwise not.
@@ 2.2.1 Predicting subcohorts in breast cancer metastasis
Recall that, the sensitivity of Algorithm 1 is the percentage of metastatic cases (out of the hold-out set of 500) that are correctly classified to be in the subcohort, while the specificity is the percentage of metastatic cases in the identified subcohort. The number of genes used by the test is the number of coefficients in the tuple $h$ whose magnitude is within a factor of 10 of the maximum magnitude coefficient in $h$. Due to well known results from compressed sensing theory, the number of genes used should be roughly proportional to the sparsity parameter input to the algorithm.
@@
One expects that as the number of genes increases, so does the specificity and the sensitivity. That is indeed the case in Figure 1. Moreover, one may increase specificity at the cost of sensitivity – simply run Algorithm 1 twice, on two different subsampled sets of patients, and obtain two tests $h_{1},h_{2}$. The new subcohort now consists of the intersection of the previously identified subchorts, i.e., both tests $h_{1},h_{2}$ have to place patients in the subcohort for the patient to be in the new subcohort. Clearly, specificity should increase for the new conservative subcohort compared to the initial subcohorts and sensitivity should decrease. Indeed that is the case – specificity increases to 70% while sensitivity decreases by a factor of 3 (see Figure 1(b)). One may even combine three or more tests in this way. Thus one can use Algorithm 1 to increase the specificity of the subcohort at the cost of decreased sensitivity, i.e., decreased size of the identified subcohort.
@@
(a) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using a single test from Algorithm 1. Sensitivity is proportional to the height of the blue bar, while Specificity is proportional to the ratio between the height of blue and red bars. The sparsity parameter (x-axis) is inversely proportional to the number of genes used.
@@
(b) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using two tests (as opposed to one), generated using two runs of Algorithm 1. Sensitivity is proportional to the height of the blue bar, while Specificity is proportional to the ratio between the height of blue and red bars. The sparsity parameter (x-axis) is inversely proportional to the number of genes used.
@@
Figure 1: Sensitivity vs Specificity trade-off using Algorithm 1, for the dataset CSC+ (12). Left: Algorithm 1 outputs a single test hyperplane that identifies subcohorts with 57% specificity, for sparsity parameter $s=6$. Right: When we use the intersection of subcohorts using two tests, the specificity increases to about 70%, but identified subcohort size (sensitivity) decreases to a third, with $s=6$. Note: the single test in the left figure uses about 15 nuclear receptors to identify subcohorts, while the two tests in the right figure use about 45 nuclear receptors.
2.2.2 Associating changes in methylation percentages with changes in nuclear receptors
A central question in genomics is to identify which specific epigenetic changes can lead to clinically significant genetic changes downstream (see for example CH (19)). In this subsection, we use our identified subcohorts to associate epigenetic changes in breast cancer related genes to changes in expression level of nuclear receptors (see Figure 3 for an example). The use of subcohorts is critical because when one averages over the entire dataset the confidence intervals become large (see for example Figure 2), and thus one can not make claims with any degree of statistical confidence.
@@ 2.2.2 Associating changes in methylation percentages with changes in nuclear receptors
Recall that, methylation of certain tumor suppressor genes can potentially lead to breast cancer progression (see SK (15)), and nuclear receptors are a target for a large number of drugs. Therefore, if we compute the mean expression levels of the two sets of genes over the entire patient population in the dataset, and find a tumor suppressor gene that is significantly more methylated and a nuclear receptor that is significantly over-expressed; then we can test whether a drug targeting that over-expressed nuclear receptor is effective and use it in new patients that show the same pattern of methylation and over-expression. Unfortunately, the situation is not so simple – there are no such sets of tumor suppressors and nuclear receptors where significant change in one is associated with significant change in the other. Figure 2 shows that the confidence intervals are just too large, when we average over the entire dataset (because of inhomogeneity in the disease). [^8] Note: Throughout this paper, the error bars are one standard deviation in length each way.
However, now that we can identify subcohorts of patients, where the identified subcohorts are described by a small set of genes; it is possible that the disease state amongst the patients in the subcohort is homogeneous enough to allow for narrow confidence intervals. That would allow us to identify any significant correlation between changes in tumor suppressor methylation and changes in nuclear receptor expression, amongst patients restricted to a subcohort. Indeed that is the case!
@@
Using Algorithm 1, and thousands of random subsamples of inputs (of 100 patients each), from the original dataset (of size 2000 patients); we generated thousands of patient subcohorts, where each subcohort had a test associated with it. The test could be either based on methylation percentage levels or nuclear receptor expression levels. Next, for every pair of subcohorts, if two subcohorts have more than a certain percentage of patients in common (we used a 10% due to limited dataset size) i.e., they look similar as far as the constituents, then we compute the mean expression and confidence interval for each gene using the common patients. If the subcohorts are made of patients in similar disease states then the confidence intervals should be narrow. And, indeed they are! In Figure 3 we compare expression levels from one such subcohort. Patients in the subcohort are identified by the test, where the coordinate weights (coefficients) for the test hyperplane are given in Figure 3(c,d). This subcohort shows significant methylation increase in BRCA2 and MSH6. Moreover, nuclear receptor NR1H2 (equivalently LXRB) is significantly over-expressed when compared to the average non-metastatic patient in the dataset. Therefore, LXR-inverse agonists (see for example CVS+ (19)) may prove to be a useful therapeutic target for this subcohort of patients, if indicated by further clinical evidence.
@@
(b) The methylation percentages in the identified subcohort, consisting mainly of metastasis cases, that are predicted to differ significantly in expression from their mean population values. In particular, for patients in this subcohort, BRCA2 and MSH6 were methylated more, both are related to DNA repair, and errors can lead to breast cancer progression SBS+ (19). Statistically insignificant bars are not shown.
@@
2.2.3 Identifying "good" subcohorts
Recall that, searching over every pair of potential subsampled inputs and subcohorts is a computationally intractable problem, since there are an exponential number of such pairs. For the small dataset with 2000 patients, and subsamples of size 1000 patients, we may need to search through $2^{1000}$ runs of Algorithm 1! However, we can train a neural net to predict when we will have subcohorts of patients that are likely to lead to statistically significant associations, i.e., the pairs of subcohorts of patients will have enough common patients to lead to narrow confidence intervals. We can then prioritize our confidence interval calculations for such subcohorts. We implemented this prioritization using a simple single hidden layer neural net. We plot the quality of the neural net model, i.e., the errors in Figure 4 in the supplement. Even with a single hidden layer neural network, we see reasonable error margins. Thus prioritization is an easy problem to learn.

### v2

@@ 3 Underlying theoretical problem
Without loss of generality, we may assume that the input also specifies a vertex $u_{0}\in S_{U}$, since one can always iterate over all vertices in $U$. Note that, $S$ is large and $|S_{U}|\gg|S_{V}|$, together imply that the cut $(S_{U},V\setminus S)$ is large. Therefore, one may formulate the above as a MAX-CUT type integer program, where one associates a $\pm 1$ variable $z_{u}$ with each of the $n$ vertices in $G$, and wants to maximize the objective:
@@
    $\langle h,x_{u^{\prime}}\rangle\geq\theta(1-z_{u^{\prime}}z_{u_{0}})$ (3)
  • Sparsity: We want the designed test to use few genes, i.e., we want its coefficients to be sparse. Hence we bound the $\ell_{1}$ norm of its coefficients (see for example rauhut) by an input parameter $s$, which is expressed as:
@@
It is worth noting that the formulated problem above is very similar to maximum margin classification using linear support vector machines. Thus, it is possible to generalize it to non-classifiers using the "kernel trick" (see for example vapnik). The only exception being the sparsity constraint above, which needs to be adapted for the chosen kernel. It may be possible to generalize the sparsity constraint for low degree polynomial kernels using standard compressed sensing methods (see for example rauhut). However, such non-linear extensions are not the focus of this work.
@@
${\forall u^{\prime}\in V:}$ $\ {\scriptstyle\langle h,x_{u^{\prime}}\rangle\geq\theta(1-\langle v_{u^{\prime}},v_{u_{0}}\rangle)}$ (8)
@@
where $x_{u}$ are the embeddings in $\mathbb{R}^{d}$, $s$ is the sparsity parameter, and $\theta$ is the separation threshold parameter, all to be specified with the input. Algorithm 1 essentially maps the vector solution of the SDP to $\pm 1$ values, to obtain the subcohort (corresponding to $\{i\in U\cup V:\ z_{i}=1\}$).
@@
1:Input: Two sets of vertices $(U,V)$ labeled 0 and $1$ respectively, embedded in $\mathbb{R}^{d}$; vertex $u_{0}\in U$, sparsity parameter $s\in(0,\infty)$, and separation threshold $\theta\in\mathbb{R}$.
@@
4:Solve the SDP relaxation (Equation 5) to obtain the solution vectors $\{v_{i}\}_{i\in[n]}$.
@@
[…]
@@
In order to show that the sensitivity and specificity of the test output by the SDP solution above are not too far from the optimum, we prove the following approximation guarantee (see the appendix, at the end, for proofs).
Theorem 3.2 (Sensitivity)
Let $S^{*}=(S^{*}_{U},S^{*}_{V})$ denote the optimal subcohort such that $|S_{U}^{*}|\ll|U|$, then Algorithm 1 computes a subcohort $S=(S_{U},S_{V})$, such that $|S_{U}|\geq\alpha|S^{*}_{U}|$, where the constant $\alpha\simeq 0.82$. Equivalently, $\frac{|S_{U}|}{|U|+|V|}\geq\alpha\cdot\frac{|S^{*}_{U}|}{|U|+|V|}$, i.e., the sensitivity of the output test is within $0.82$ of the optimum.
Theorem 3.3 (Specificity)
For $\kappa\in(0,1)$, let $\kappa\cdot|U||V|$ denote the fraction of edges in the optimal sparse cut separating $U$ from $V$. The specificity of the test output by Algorithm 1 is lower bounded by the solution to the following optimization problem:
$\min_{x,y\in(0,1)}\frac{x\cdot|U|}{(1-y)|V|+x|U|}\qquad\mathrm{s.t.}\quad x\cdot y\geq\tilde{\alpha}\cdot\kappa,$ (9)
where $\tilde{\alpha}=\frac{\alpha}{2}\simeq 0.41$.
Note that the solution of the optimization problem in Equation 9 grows as $\Omega(\sqrt{\kappa})$, when $|U|$ is assumed equal to $|V|$ (this can be seen by substituting $y=\frac{\tilde{\alpha}\cdot\kappa}{x}$ in the objective). A plot of the solution surface, together with the proof, is given in the appendix (Section 7).
While the optimum hyperplane separator $h$ remains unchanged after rounding though the input specified threshold $\theta$ is perturbed during rounding. However, we may bound the perturbation amount by using Grothendieck’s identity, i.e.,
@@
$\langle h,x_{u^{\prime}}\rangle$ $\geq$ $\theta\left(1-\frac{2}{\pi}(z_{u^{\prime}}z_{u_{0}})\right)$ (12)
@@ 4 Experiments: Genomics applications
  • In Subsection 4.1, we examine the quality of the identified subcohorts. A random test, i.e., a random half-space, can be expected to generate subcohorts with less than 50% metastasis cases (as the number of non-metastasis patients is higher in the dataset). In Figure 1(a), we show that the linear tests generated by Algorithm 1 identify subcohorts where greater than 50% of the cases are metastatic (thus outperforming random). Moreover, if instead of using one hyperplane to identify a subcohort, we use two hyperplanes to identify a subcohort then the identified subcohorts have 60-70% of their cases as metastatic (see Figure 1(b)). We also compare our algorithm’s specificity (for a given sensitivity) with a version of the well-known PRIM algorithm for subgroup discovery (this comparison is in the Appendix Section 9).
@@
  • In Subsection 4.2, we use Algorithm 1 to identify a clinically interesting subcohort of breast cancer patients – those that may respond to LXR inverse agonists. In the process, we also associate epigenetic changes with changes in nuclear receptor expressions. It is known brcgenes, that about $40$ genes act as tumor suppressors, and their methylation leads to reduced expression and potentially cancer progression. Moreover, in nrdrugs; nrdrugs2, the authors discuss the role of nuclear receptors as drug targets. Therefore, it is natural to look for associations between significant changes in nuclear receptor expressions and significant changes in methylation levels of breast cancer related genes (over population baseline levels) – the latter can guide treatment drug targets from the former. We show that, even though at the entire dataset level such associations do not exist, but at the subcohort level they do exist (see Figure 3).
@@
Dataset: The results in this section are based on the METABRIC dataset metabric; metabric2; metabric3, which contains RNA microarray data from biopsy samples of approximately 1904 (911 with positive lymph nodes and 993 without positive lymph nodes) breast cancer patients, and methylation data for approximately 13K genes in about 1406 (669 with positive lymph nodes and 737 without positive lymph nodes). In this paper, we (crudely) denote patients with one or more positive lymph nodes upon biopsy as metastasis cases, while the rest as non-metastasis cases.
4.1 Predicting breast cancer metastasis
@@
Recall that, the sensitivity of Algorithm 1 is the percentage of metastatic cases that are correctly classified to be in the subcohort, while the specificity is the percentage of metastatic cases in the identified subcohort. The number of genes used by the test is defined as the number of coefficients of $h$ whose magnitude is at least $\frac{1}{10}$ of the maximum magnitude coefficient in $h$. Due to well known results from $\ell_{1}$ norm minimization rauhut, the number of genes used should be roughly proportional to the sparsity parameter input to the algorithm (Equation 4).
Figure 1 shows the sensitivity, specificity and sparsity trade-offs for our subcohort identification algorithm. To generate Figure 1, Algorithm 1 uses the nuclear receptor expressions of a subset of 50 metastatic and 50 non-metastatic cases from the full dataset, the sparsity parameter $s$, a threshold $\theta$, and an arbitrary metastasis patient that is assumed to belong to the subcohort to be identified. The algorithm outputs a test $h$, where $h$ is just a tuple of weights for the nuclear receptors. On a separate hold-out set of a 1000 patients (500 with metastasis and 500 without), the test simply calculates the weighted average value of the given nuclear receptor expression levels for each patient in the hold-out, and if this value is below the threshold then the patient is added to the subcohort, otherwise not. The above exercise is repeated $10$ times for various sparsity values, and sensitivity and specificity measured each time, to arrive at Figure 1.
@@
Interpretation of Figure 1: First, as the number of genes increases, so do specificity and sensitivity in Figure 1. Second, sensitivity can be traded-off for specificity: We run Algorithm 1 twice, on two different subsampled sets of patients, and obtain two half-spaces/tests $h_{1},h_{2}$ that correspond to two subcohorts. The new subcohort is defined to be the intersection of the previously identified half-spaces, i.e., both tests $h_{1},h_{2}$ have to place patients in the subcohort for the patient to be in the new subcohort. Clearly, specificity should increase for this more conservatively defined subcohort and sensitivity should decrease. Indeed that is the case – specificity increases to 70% while sensitivity decreases by a factor of 3 (see Figure 1(b)). One may even combine three or more tests in this way. Thus one can use Algorithm 1 to increase the specificity of the subcohort at the cost of decreased sensitivity, i.e., decreased size of the identified subcohort.
Comparison: We compare our algorithm with a version of the well-known PRIM algorithm in Appendix (Section 9). It is clear from our experiments that the specificity of the MAX-CUT based approach outperforms that of the local search based heuristic when sensitivity is held constant. This is despite the fact that the local search heuristic was allowed an order of magnitude more time for computations.
@@
(a) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using a single test from Algorithm 1. Sensitivity is proportional to the height of the blue bar, while Specificity is proportional to the ratio between the height of blue and red bars. The sparsity parameter (x-axis) is proportional to the number of genes used.
@@
(b) The number of metastasis cases (true positives) and non-metastasis cases (false positives) in the identified subchort using two tests (vs one in L figure), generated using two runs of Algorithm 1. Sensitivity is proportional to the height of the blue bar, while Specificity is proportional to the ratio between the height of blue and red bars. The sparsity parameter is proportional to the number of genes used.
@@
Figure 1: Sensitivity vs Specificity trade-off using Algorithm 1, for the dataset metabric. Left: Algorithm 1 outputs a single test hyperplane that identifies subcohorts with 57% specificity, for sparsity parameter $s=6$. Right: When we use the intersection of subcohorts using two tests, the specificity increases to about 70%, but identified subcohort size (sensitivity) decreases to a third, with $s=6$. Note: the single test in the left figure uses about 15 nuclear receptors to identify subcohorts, while the two tests in the right figure use about 45 nuclear receptors.
4.2 Methylation and gene expression
Algorithm 2 Identify Critical Genes
1:Input: Dataset $D\subset\mathbb{R}^{d}$, and parameter $\epsilon\in(0,1)$ (the $d$ coordinates are the genes)
2:Output: A subcohort with critical genes that are significantly differently expressed from the population
3:
4:Uniformly sample $S\subset D$ with $|S|\simeq\epsilon\cdot|D|$
5:Use Algorithm 1 on $S$ to compute a half-space $h$ corresponding to a likely homogeneous subcohort.
6:Compute means and standard deviation of each of $d$ coordinates in population and subcohort.
7:If any coordinate has mean which is one standard deviation away from population mean, then return the coordinate (gene), subcohort and $h$ (test)
8:Else generate a new sample (Step 1) and repeat
One application of Algorithm 1 is to identify genes which differ significantly between a subcohort and the population. Algorithm 2 presents the basic idea. We will use a slight variation of it to relate epigenetic changes (nucleotide methylation changes) to clinically significant genetic changes downstream – an important question in genomics (see for example epi1). The use of subsets is critical as averaging over the entire dataset leads to large confidence intervals.
@@ 4.2 Methylation and gene expression
Background: Methylation of certain tumor suppressor genes can potentially lead to breast cancer progression epi1; brcgenes, and nuclear receptors are a target for a large number of drugs nrdrugs; nrdrugs2. Therefore, if we compute the mean methylation and expression levels of critical genes over the entire patient population in the dataset to find a tumor suppressor gene that is significantly more methylated and a nuclear receptor that is significantly over-expressed; then we can test whether a drug targeting that over-expressed nuclear receptor is effective in patients that show a similar pattern of excessive methylation and over-expression. Unfortunately, the situation is not so simple – there are no such sets of tumor suppressors and nuclear receptors where significant change in one is associated with significant change in the other. Figure 2 shows that the confidence intervals are just too large, when we average over the entire dataset (because of inhomogeneity in the disease). However, it is possible that the disease state amongst the patients in some subcohorts is homogeneous enough to allow for narrow confidence intervals, and thus identify any simultaneous significant deviation in tumor suppressor methylation and nuclear receptor expression, amongst patients restricted to such a subcohort. Indeed that is the case below.
Algorithm 2 for relating epigenetics with nuclear receptors: We use a variant of Algorithm 2, with about $1000$ different random subsamples of inputs (of $100$ patients each), from the original dataset (of size $2000$ patients) to generate patient subcohorts of breast cancer patients (and corresponding tests to identify the subcohorts) in two ways: (1) based on methylation percentage levels, and (2) based on nuclear receptor expression levels.
Next, for every pair of subcohorts generated (one from methylation levels and another from nuclear receptor expressions), if the two subcohorts have more than a certain percentage of patients in common (we used a threshold of 10% due to limited dataset size) i.e., they share a relatively large subset of the patient population, then they are likely to be homogeneous as two different sets of markers – genetic and epigenetic – are identifying a similar subset of patients.
Next, we compute the mean methylation and expression together with confidence intervals, for each gene using the common patients. Finally, we iterate until we find a patient subcohort with genes that have large deviations from the population means.
One such subcohort is shown in Figure 3.
@@
Interpretation of Figure 3: Patients in the subcohort are identified by sparse linear tests, where the coordinate weights (coefficients) for the test hyperplane are given in Figure 3(c,d). Interestingly, this particular subcohort of patients shows significant methylation increase in BRCA2 and MSH6. Moreover, nuclear receptor NR1H2 (equivalently LXRB) is significantly over-expressed when compared to the average non-metastatic patient in the dataset. Therefore, LXR-inverse agonists (see for example lxr-ant) may prove to be a useful therapeutic target for this subcohort of patients, if indicated by further clinical evidence.
@@
(b) The methylation percentages in the identified subcohort, consisting mainly of metastasis cases, that are predicted to differ significantly in expression from their mean population values. In particular, for patients in this subcohort, BRCA2 and MSH6 were methylated more, both are related to DNA repair, and errors can lead to breast cancer progression brca2. Statistically insignificant bars are not shown.